On November 4th, against the Houston Rockets, Ron Artest scored a total of 15 points, making a total of 9 shots. He scored all of his points on 2-point and 3-point shots. How many of each type did he make?
Answer by richwmiller(17219)

a+b+c=9,
2a+3b+0*c=15
All 3 pointers 5 shots 3*5
All 2 pointers 7.5
There is no way there were 9 good shots with a total of 15
So he must have missed some
1*3+2*6
1 3 three pointer and 6 2 pointers and and 2 missed shots
1+6+2=9
2 no
3 no
4 no
5 three pointers no 2 pointers and 4 missed